OpenStudy (vanz):

whats an real world example for open sentences? like i would use open sentences if i had to...

myininaya (myininaya):

I'm not sure what an open sentence is. This sounds like an english question. Could you try posting this in the english section. Thanks. And sorry for that spammer.

OpenStudy (vanz):

No this is math.

myininaya (myininaya):

What is an open sentence?

myininaya (myininaya):

lol and sorry i didn't know it was just never heard of it

myininaya (myininaya):

Examples of open sentences include: 3x − 9 = 21, whose only solution for x is 10; 4x + 3 > 9, whose solutions for x are all numbers greater than 3/2; x + y = 0, whose solutions for x and y are all pairs of numbers that are additive inverses; 3x + 9 = 3(x + 3), whose solutions for x are all numbers. 3x + 9 = 3(x + 4), which has no solution. There are some examples I found of open sentences.

myininaya (myininaya):

let me see if i can look at the pattern and determine a real world example

OpenStudy (vanz):

Oh no its okay now! I got the answer haha.

myininaya (myininaya):

We need to put some money into our bank account such that are account is no longer negative. We could put in 25 dollars or more in order for our balance to be 0 or positive.

myininaya (myininaya):

Hey vanz thanks for giving me vocabulary.

myininaya (myininaya):

more vocabulary*

OpenStudy (vanz):

Do you have anything for equations of a line? and no problem thanks for the help. :)

myininaya (myininaya):

y=mx+b anything that can be written in this form is a line

myininaya (myininaya):

or even x=constant or y=constant

myininaya (myininaya):

these are all lines and if an equation can not be written in any of these forms then it is not linear is that what you were looking for?

OpenStudy (vanz):

I am looking for real world examples like when would I ever use this

myininaya (myininaya):

oh

myininaya (myininaya):

hmmm.... I know you can use it some jobs that involve technology.

myininaya (myininaya):

I cannot think of an example off the top of my head.

myininaya (myininaya):

let me see if i can find someone who can think of an example

OpenStudy (vanz):

okay Thank you! I'll wait.

myininaya (myininaya):

line goes on forever what can i say goes on forever?

myininaya (myininaya):

i guess we could say that we can make two linear equations for two train tracks whose tracks are almost linear and find their approximate intersection

OpenStudy (vanz):

That works! :D

myininaya (myininaya):

thats the best i can think of lol

OpenStudy (vanz):

do you have any examples for system of linear? :D

myininaya (myininaya):

that was it actually

myininaya (myininaya):

that was a system of linear equations above

OpenStudy (vanz):

like i can use it as an equation of a line example AND linear equations?

myininaya (myininaya):

i don't see why not